A car covers distance of 2 km in 2.5 min.if it covers half of the distance with speed 40km/hr the rest distance t shall cover with a speed of?

Grade:12th pass

1 Answers

Khimraj
5 years ago
time taken by car to cover half distance with velocity 40km/hr
t = 1/40 = 1/20 hr
remaing time = 2.5/60 โ€“ 1/20 = 1/24 โ€“ 1/40 = 1/60 hr
velocity required to cover 1 km in 1/60 hr
v = 1/(1/60) = 60km/hr
Hope it clears.ย